If I have 80 feet of fencing and I need to have 500 square feet of area. How would I do it?
Answer by JulietG(1812)

Your perimeter is 2Length + 2Width. You know you have 80 feet of fencing, so 2L + 2W = 80. Divide each side by 2 to get L + W = 40
.
Width * Length = area. W * L = 500
.
What you need is 2 numbers that equal 40 when added and 500 when multiplied.
What are the factors of 500?
1,500
2,250
4,125
5,100
10,50
20,25
None of these work. Are you sure the 80 and 500 numbers are correct/